OpenStudy (anonymous):

How to find GCF of 27 and 45, when I already have 3 cubed and 3 squared under it and 5??? Please help

OpenStudy (anonymous):

27 = 3*3*3 45 = 3*3*5 the gcf is greatest number which divides into 27 and 45 take out one of each common factors that is 3 and 3 GCF = 9

OpenStudy (anonymous):

would you like to try one yourself - how about 16 and 24?

OpenStudy (anonymous):

Is the answer 4?

OpenStudy (anonymous):

16 = 2*2*2*2 24 = 2*2*2*3

OpenStudy (anonymous):

gotta go now

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Thank you, goodbye

OpenStudy (anonymous):

yup yw